UT Feb 14 2022 - DEC Night (Full)



Staff:

  • OSA: J.Williams
  • ISp: Mtn
  • Mtn Manager: B.Waggoner
  • ISA/Observer: O.Kuhn, with C.Veillet, H. Choi for TMS, and with B. Rothberg for LUCI1-AO
  • AO: J.C. Guerra, D. Miller
  • Special Guests:

Night Info:

  • Sunset 18:04
  • Nautical Twilight Ends 18:58
  • Astro Twilight Ends 19:26
  • Night Center 00:33
  • Astro Twilight Begins 05:41
  • Nautical Twilight Begins 06:09
  • Moonset 06:11 - Moon 93% Illumination
  • Sunrise 07:03

Night Plan:

  • Starting Configuration LBCs

LBC/TMS Testing

  • LBC/TMS Testing twilight-23:30
    • Active/Passive mode (Blue active and Red passive)
    • 3 fields to go up to 23:30 while staying away from the Moon. Sidereal tracking. 1mn exp time. g/r
      • 02 32 50 +00 18 30 (2021 VK27) - up to ~21:00

      • 03 02 20 +14 09 50 (2020 VL26) - up to ~22:00

      • 04 00 05 +39 52 15 (2015 RG328) - up to 23:30

Contingeny - if LBC/TMS cannot operate

  • Decision by 15hr daytime
  • LBTI CQ Tau w/ALES+vAPP coronagraph
    • photometric conditions
    • Seeing < 1.2"
    • 4hrs
  • Reconfigure for LUCIs - 23:30-00:00

LUCI-1 AO Testing

  • LUCI-1 AO Testing 00:01-twlight
    • OT files Location: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
      • ESM_V34.xml
      • AO_Testing_V64.xml
    • OSCO
    • Bin 3 Testing GOpt + NCPA
      • On Axis stars from folder AO STARS - FULL DITHERS NEW in OT
      • Select stars based on RA, elevation, and brightness as needed
    • Repeat the UVa Jets Bin 3 program to replicate issue seen - only need some exposures not full script execution
      • Remote working directory: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM
      • OT file: S255_SOUL_NIRJets_OT.xml
      • object sets at 08:30 UT
    • Bin 1 Testing
      • On Axis stars from folder AO STARS - FULL DITHERS NEW in OT
      • Select stars based on RA, elevation, and brightness as needed
    • ESM Bin 1
      • M61 - face on spiral Bin 1 (R=10.1) - Ks-band - midnight-twilight
      • M92 - Bin 1 03:15hr-twilight - provides PSF characterization
      • Backup BIn 1 targets
        • M61 - face-on spiral (R=10.1) Ks-band - handover to twilight - top priority
        • M102 - edge on spiral Bin 1 (R=10.9) H-band - - 01hr-twilight - only if M61 can't be done
        • NGC 2392 - Bin 1 (R=8.0) - compact Nebula twilght-02:30 - backup target if time - can skip the P-beta filter
    • ESM Bin 2
      • M67 - Bin 2 handover-04hr - PSF characterization - first priority for this bin
      • Dual BLR target (Bin 1 or BIn 2 - Ref star is borderline) 03:30hr-twilight - medium priority
    • Confirm Bin 2 OK (if no Bin 2 run during OSURC run) - medium priority
      • On Axis stars from folder AO STARS - FULL DITHERS NEW in OT
      • Select stars based on RA, elevation, and brightness as needed
    • Confirm Bin 4 OK - medium priority
      • On Axis stars from folder AO STARS - FULL DITHERS NEW in OT
      • Select stars based on RA, elevation, and brightness as needed
    • Pretty Pix for Full AO - lower priority
      • Frosty Leo - handover - 05hr
        • On-Axis - R=9.16 (tricky to see if star is actually visible)

Poor Weather Backup

  • Local Mergers - Rothberg
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V40.xml
    • Any conditions (guiders up to 4" - as long as guider can lock on and WFS) - no particular order
      • UGC 9829 - 2nd half
      • UGC 10607 -2nd half
      • UGC 10675 -2nd half
      • NGC 4004 -2nd half
      • Arp 156 -2nd half
      • NGC 3597 -00:00-03:30hr
    • If clear and poor-poor seeing (track mode if seeing so poor guider loses lock)
      • LUCI-Bino Standards
        • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml
        • If seeing > 4" - make these track modes in the OT and export

Log

LBC/TMS (Blue Active/ Red Passive to start, and then Red Active/Blue Passive for the remaining time)

01:01 Josh is opening the enclosure

01:06 Heejoo is rebooting the TMS system

01:07 Sunset

01:12 TMS is ready - the sky is a bit too bright still.

01:20 No old reference - the TMS system has been realigned.

01:25 Slewing to nearby field Blank 02+13 for g & r skyflats (scale should be closer to 2:1 for g:r)

01:58 12-deg twilight

2021 VK27 (blue-active and red-passive)

01:52 Slewing to the copointing field for 2021 VK27

01:58 Running dohybrid (no earlier reference - stars in sky flats have a lot of coma)
  • face-back is 0.083 SX and 0.031 DX currently
  • blue converged after WRS only, but bounced back up while red was converging.
02:06 Repeating just dofpia

02:12 Setting the reference. lbc_set_tms_offset.py -s 2

02:15 ./tms_loop_left.pl

02:16 Started the science script.
  • The first images look good - FWHM 3.0 on blue and 2.8 on red (pixl)
02:23 Changed the exptime to 40 sec to prevent active optics Z4 corrections. Earlier there was one (-529nm) sent (and presumably backed out by TMS) on the red side, but none on the blue side. Now no Z04 corrections are being sent from the focus chips.
  • elevation 45 deg
  • last blue image 3.3 pix and last red image 3.1 pix
02:56 dofpia,/redonly

Defoc   AstX   AstY   ComaX  ComaY   Sph        |  Est.   Est. |         
 (Z4)    (Z5)   (Z6)   (Z7)   (Z8)   (Z11)  (Z22)| seeing  WIQ  |         
--------------------------------------------------------------------------
  -694    100    131    235   -120   -357    -51   0.42   0.39     red

03:00 Starting the science script - still on 2021 VK27
  • 03:14 blue IQ is 4.1 and red is 2.6 pix - elevation 37 deg.

2020 VL26 (red-active and blue-passive)

03:14 Slewing to copointing field for 2020 VL26.

03:17 Control-C TMS loop and run dofpia on both sides.

03:23 Send copointing script and set a new reference here.

03:25 Starting the tms loop on the right side

03:26 Sending copointing corrections.

BLUE 4.3" and RED 10.6" from rotator centers of LBCB and LBCR, respectively
COPOINTING: B=32430 R=32427
Pointing updates: delta_IE = 3.44", delta_CA = -6.13"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.09    -0.13    -2.76    -1.92
              DX:   -0.09     0.16     3.41     2.00

03:27 Starting the science script (using 40-sec) exposures and with tms active on the DX side.
  • elevation 48 deg
  • FWHM on first blue image 2.8 pix and on the first red one, 2.6 pix
04:01 Starting the second round
  • FWHM on blue now ~3.3 - there is an aberration developing. face-back -0.130 now.
  • FWHM on red ~2.7 pix - looks like the collimation is holding. face-back 0.017 now.
  • Mirrors: SX: tglass=3.031, tamb=2.82 and DX: tglass=2.814, tamb=2.956
  • Blue IQ is definitely degrading - FWHM 4.3 and aberrated.
    • Blue face-back -0.122 now (vs red face-back 0.021).
  • 04:30 Stopping to recollimate the blue side
  • 04:30 Slewed to the copointing field
    • dofpia, /blueonly while keeping the tms loop running on the right side.
      • 5 iterations were needed, net Z4 was -3000 nm
  • 04:38 Resumed science script
    • blue IQ is better, though now elongated 90 deg from before. Red IQ is holding ~ 2.5 pix
  • 04:46 red IQ ~2.6 pix and elevation 33 deg
  • 04:55 stopping at elevation ~31 deg

2015 RG328 (red active, blue passive)

04:56 Slewing to the copointing field with TMS DX side only running
  • 50 deg elevation
04:58 dofpia, /blueonly
  • 4 iterations were required. Net z4 was about 800 nm

05:03 copointing

-------------------------------------------------------------------------
COPOINTING: B=50443 R=50426
Pointing updates: delta_IE = 4.49", delta_CA = 6.25"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
             SX:    0.01     0.04     0.80    -0.14
              DX:   -0.01    -0.00    -0.01     0.27
-------------------------------------------------------------------------

05:06 Starting the science (40-s)
  • First images have FWHM 2.5 (red) and ~3.1 (blue)
  • lbcr 050827 has a small jump. This is the 2nd image of the sequence.
  • lbcr 051204 also shows a small jump.
  • the next one, lbcr 051316 looks nice and round, FWHM 2.55 pix.
  • lbcr 051429 also seems a bit ew elongated - blue one 051400 looks ok
  • 05:30 FWHM on blue 3.7 and on red ~2.7 - blue seems to be developing an aberration
05:42 Starting the second round on the science
  • FWHM on blue 4 pix and on red ~2.7 pix --- elevation 41 deg
    • SX face-back 0.0 deg, DX face-back -0.059 deg
  • 05:57 red IQ still ~2.6 pix
  • 06:11 but the blue side is developing an aberration

06:19 Slewing back to the copointing field and allowing a group rotation update to DX before control-C'ing TMS

06:21 runing dofpia on both sides

06:23 Started ./tms_loop_passive.pl during dofpia
  • 5 iterations were required and corrections: z4, z11 and z22 were sent on LBCR. Only Z4 on LBCB.

06:36 Control-C'd the passive loop.

LUCI1-AO

06:36 Josh is reconfiguring to LUCI

07:01 Checking pointing

OSCO

07:08 Slewing to OSCO star HD265589
  • R= 7.52, OCAM said 8.6
  • Failed on center pupils
  • Bin 1, NCPA on
  • Loop is closed, but the script is hung and telsvc gui reports "Wait for Collimation..."
    • On LUCI Observation Execution Panel, click Other... --> Abort and skip to the next line
    • luci1.20220214.0004.fits is the first exposure.
    • FWHM 2.93, 3.22 on Aladin display
  • 07:30 Finished

S255 script Bin 3 (are the jumps seen last night repeatable? IT)

07:33 Slewing to S255 --- bin 3, low elevation (where we saw a jump last night).
  • OCAM says the star has R=14.8, same as last night (bin3).
  • When the loop was closed, the script continued
  • L1: luci1.20220213.21-22 object & sky at Ks with NCPA off
    • FWHM ~ 7 x 4-5
    • elev 37 deg
  • L1: luci1.20220213.23- NCPA on
    • the elongation is of the same sense as before but a bit better.
    • frames near the end of the cube 25 show a strong elongation.
    • all frames in cube 28 look elongated in the same sense
    • before 30, image in CCD47 became elongated - jumped - came back and them took time to settle to round again.
08:03 Bin3, higher elevation
  • S255 was setting, trying Bin 3 at higher elevation
  • L1 33-36
  • Slewing to Star3_bin3_R14.9_PA0
  • We are seeing an error on GCS and not getting new images. getting new blocked sending - closed to edge on GCS. WFS images are not coming in. Resending the preset. OK now.
  • OSCO says 16.3
  • All of the frames of the cube 33 show two stars.
  • In CCD47 - it was hard to see at first, but with the loop closed it's clear this is a double star, with one component brighter than the other. This will affect performance.
08:26 Bin 3, higher elevation, GD296 (hoping it's not a double)
  • No star found in guider camera FOV. Josh is checking pointing, but no correction was needed. We sent the preset again and the star was not found. bin3, R=15.3, PA=0
08:35 Same star as before, GD296, but different PA, PA=120.
  • Closing the loop failed on center star - there is no star on CCD47.
    • Image 37 - is the star in LUCI FOV? There is a star at 1088, 189. and the whole field shifted.
    • Mystery --- at PA=0, star was not in GCS but was on CCD47. At PA=120, star in GCS but not on CCD47.

Characterizing the effect of moonlight

08:43 Moon angle test, 11 deg from the moon
  • Ton899, R=14.37 and 11.1 deg from the moon, OCAM says magnitude 15.4
  • Closing loop failed on preset AO (AO preset was not cancelled from FLAO GUI)
  • Trying again - center pupils failed.
  • Again - the loop is closed, but the estimation of the pupils is wrong - dimension of pupils wrong. Camera lens is mid-range.
08:56 Moon angle test, 14.5 deg from the moon.
  • 2MASS J09081129+2434084, R=10.91 (OSCO says 12.2)
  • I sent the preset, but we are not slewing. telsvc says "Wait for Collimation...". The preset has been cancelled, I tried to skip "wait for collimation." and again to reconnect to IIF. I also tried to abort the script (although no script was running). Around the time we reauthorized, we saw a message go by about the 1200-sec timeout - that aborted the script.
  • Josh reauthorized finally and that helped. Sounds like IT 5008.
  • I sent the script again - no guide star found. Pointing check done.
  • 09:13 I sent the script again
    • Closing the loop. OSCO says mag 12.2
    • 09:26 Accidentally sent the preset again. Stop AO - loop opened.
    • subwin 256 x 256 (SW 1 923 1018 256 256)
    • 09:27 starting 46- 50 --- these are with NCPA off
      • 46 shows a jump
    • Strehl 61%
    • 09:41 turning NCPA on for image 51 and onwards
    • Images with NCPA off look better.
    • 58 shows a jump - return from a large offset, but the jump was near the end of the 100 DIT cube.

Pretty Pic

09:57 Frosty Leo on-axis
  • R=10.4 but nothing seen on CCD47 (the star could be obscured) - cancelling this.

ESM

Bin 1 ESM
10:01 M61 Bin1 Bino ESM & SL
  • 10:18 Sending the preset: TTM Active left and Active right
  • Star not seen very well in CCD47 -
10:32 M102 Bin1 BinoESM & SL
  • We had an issue with a duplicate left preset - though the preset had been canceled (is this like IT 8534?)
  • WFS hit a limit - we have to change the rotator angle, but the re-rotate is still out of range.
10:35 M92 R9pt4_ESM_SL (Bin 1 ESM)
  • re-rotator is at 262 but out of range. This script worked before - same time of night, same rotator position for LUCI.
  • loop is closed (LUCI preset failed, however).
  • Aborting and canc
10:47 Trying again M92 R9pt4 ESM_SL (Bin 1 ESM)
  • Closing loop failed on center pupils
  • Nothing seen on OCAM viewer - Juan Carlos reconfigured(?) OCAM - even at 200 Hz he doesn't see anything.
Bin 4 ESM
10:58 Trying now the Bin 4 ESM+SL observation of M92.
  • No preset coming through. It says "Wait for Coll..." on the left side. There is nothing to clear & nothing to abort.Skip wait for collimation, Reconnnect IIF did nothing. (04:01 - Josh is reauthorizing (but not going from one instrument to another) - Barry adding a note to IT 5008).
  • 11:05 - Josh reauthorized by going to LBC and back to LUCI. Now "Wait for collim..." is gone from telsvc GUI - instead "Error" on left side. Message flashed through LUCI msg stream about 1200s timeout - the is the time before LUCI script aborts.
  • 11:06(?) resent the preset.
  • 11:14 the loop is closed - bin 4, R = 15.17, OCAM sees 16.5
  • 11:26 LUCI script starting
    • L1 62-76 and L2 1-15 --- FWHM on L1 is ~3 pix and the L2 images are really ugly - after the script finished, I initialized L2 since I wasn't 100% sure I did so earlier and then took another image, 16. The images on it looked like those on 15.
    • After L1 66 and L2 5, the script paused - but there was no pause
Bin1 ESM again (M102 again, but forced to bin2)
12:00 M102 - bright (bin1) star forced to bin2.
  • 12:05 - sending the preset
  • Now the star is seen. OCAM thinks 12.6
  • 12:08 Starting, L1: 77 and L2: 17-
  • L1 PSF is a bit double-peaked (80), L2 PSF also - see 21. - These were H band
  • Put Ks in by hand.
  • The active optics probe failed to move to the offest positoin - the AO loop paused (the AO ref star was also close to edge of AO patrol field so AO loop paused)
  • Made an absolute offset to 0 0 (sync since large). Now taking test, Ks, images: L1 84 and L2 23.
  • From L1 85 and L2 24 - but loop paused and guider out of patrol field.
  • Problem: some offsets move guide probe out of off-axis patrol field and when that happens, even when the AO ref star is in the AO patrol field, then the loop pauses.
12:35 M102 -
  • Barry used the original version of the script - as bin 1, but sent as bin 2.
  • Rmag 12.52 but OSCO says 12.50
  • FWHM on L1 is very good - 2.2 pix on 91, for example. L2 30 has a out-of-round shape
  • L1 83- and L2 - 23-
  • 12:56 L2 guider FWHM 0.40"
  • 13:15 --- Again, the LUCI execution panel just paused. The readout boxes were greyed out, but not the offset.
13:28 Juan Carlos is doing a test

13:31 Closing up.

Issues:

RB_Science crashed:

---- receiving 'lbcr.20220214.035813.fits' [frame 2] (5.0s/75.5s)
      object:2020 VL26 | filter:r-SLOAN | exptime:40.24 | pa_pnt:-0.02 | dithoffx:0 | dithoffy:0
double free or corruption (!prev)

Unknown error! Please restart (and report the following messages).
------------------------------------------------------------

20220214 directory not in Tucson repository

Tonight's data are going to the mountain repository, but not to Tucson. There is no 20220214 directory in Tucson.

-- BarryRothberg - 10 Feb 2022
Topic revision: r11 - 14 Feb 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback